### v1

# Distributed implementation of tensor-product selected configuration interaction

Abstract
In recent years, the hybrid “QC+HPC” strategy—where quantum computers screen important determinants, followed by exact diagonalization on classical computers—has shown great potential in the study of strongly correlated systems in quantum chemistry. Last year, an IBM team proposed a novel scheme that utilizes quantum computers to select important bit strings that are then used to construct a spin-adapted determinant space via tensor products. Inspired by this, we have specifically designed a completely new algorithm for this tensor-product selected configuration interaction (SCI). Moreover, for the first time worldwide, we have implemented distributed storage of the CI vector in an SCI program, enabling efficient handling of large-scale computation. Since this study is independent and does not involve determinant selection by quantum computers, we employed our SCI program to conduct full configuration interaction (FCI) computations. Our FCI calculations for N2 (aug-cc-pVDZ) under D2h symmetry and CN (cc-pVTZ) under C2v symmetry, involving $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants, respectively, exceed the previous record of $2\times 10^{9}$ determinants computed with the DICE program [J. Chem. Phys. 149, 214110 (2018)] by more than two orders of magnitude. These results set a new benchmark for SCI computations and lay the groundwork for further advancements in SCI methods.
1 Introduction
The Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry. 1, 2, 3, 4 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets. By the use of a distributed CI vector, the largest system calculated to date is propane with the STO-3G basis set, involving up to $1.31\times 10^{13}$ determinants 4.
In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 5 is known as the gold standard for weakly correlated systems owing to its balanced accuracy and efficiency. While achieving accurate calculations for strongly correlated systems remains the most formidable challenge in quantum chemistry, the selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue. As early as 1973, Huron et al. introduced the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI) method 6 in which candidate determinants are selected by rigorously estimating their energy contributions via second-order perturbation theory. Although this selection scheme is considered reliable, its high computational cost precludes its application to larger molecules. Subsequently, methods such as Adaptive Sampling CI (ASCI) 7, 8, and Semistochastic Heat-Bath CI (SHCI) 9, 10, 11 were developed, which can essentially be regarded as approximations to CIPSI.
In recent years, with the rise of quantum computing (QC), several research groups have attempted to rapidly screen important determinants using quantum devices. However, the noise inherent in quantum computers can lead to the loss of some determinants, preventing the generation of wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator. To address this issue, last year a team from IBM proposed an innovative strategy: using quantum computers to select important bit strings, and then constructing the complete determinants via tensor products. 12 In this paper, we refer to this variant of SCI,which is characterized by this structure, as Tensor-Product SCI (TPSCI).
Since CI calculations are variational, including more determinants generally improves accuracy; however, the associated memory and computational costs for diagonalization increase dramatically. Currently, the widely used DICE program, 9, 10 which distributes the nonzero Hamiltonian matrix elements across nodes, enables calculations with up to $2\times 10^{9}$ determinants. Nevertheless, DICE employs replicated storage for the CI vector, so that the memory available on a single node becomes a bottleneck, thereby limiting the maximum feasible CI vector size despite multi-node execution.
In this work, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, overcoming single-node memory limitation and enabling the use of the entire supercomputer’s available memory for larger-scale calculations. Moreover, we have designed a completely new, highly efficient algorithm specifically for TPSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku. We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 13 and in line with current trends in correlated methods. 14 The combination of low memory footprint and low precision operations make our work promising for application to leadership class supercomputers based on GPUs.
Ultimately, we computed the FCI energies of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), which, under D2h and C2v symmetry, respectively, involve $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants. These results significantly exceed the previously reported maximum SCI scale of $2\times 10^{9}$ determinants computed with the DICE program. 11 These achievements lay a solid foundation for the future development and application of TPSCI. Meanwhile, as previous FCI calculations for N2 15 and CN 16 were performed with the smaller cc-pVDZ basis set, our results serve as valuable benchmark data for further research.
2 CI Vector Distributed Algorithm for TPSCI
In this section, we will describe the parallel algorithm we have designed for TPSCI, including the important optimizations required for large-scale application.
2.1 Distributed CI Vector Storage
In the framework proposed by IBM, only important bit strings are selected, which means that the chosen $\alpha$ and $\beta$ bit strings are identical. The determinant space constructed via tensor products can thus be used to generate closed-shell singlet wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator.
In our program, the selected $\alpha$ and $\beta$ bit strings are allowed to differ. Suppose their sets are defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,x\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,y\}$, where $S^{\alpha}$ and $S^{\beta}$ are the sets of selected $\alpha$-spin and $\beta$-spin bit strings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$. The determinant space constructed via tensor products is then given by:
$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,x;\,u=1,\dots,y\}.$ (1)
The ordering of $\{S^{\alpha}\}$ and $\{S^{\beta}\}$ can be freely chosen. However, once determined, the index order of the determinants in the determinant space is also fixed. Specifically, the $K$-th determinant is denoted as $|D_{K}\rangle$ and given by:
$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{y}+1,\quad u=\text{MOD}(K,y).$ (2)
When the number of determinants exceeds the storage capacity of a single node, the CI vector must be distributed across nodes. In this work, the term CI vector refers to the array of CI coefficients stored in the same order as the determinant indices. As illustrated in Fig. 1, we distribute the determinants based on their $\alpha$ bit strings while ensuring an approximately even allocation, a strategy known as memory balance.
Since the determinants are indexed such that $\alpha$ bit strings vary first, followed by $\beta$ bit strings, we define a segment as the subset of the CI vector corresponding to all determinants sharing the same $\alpha$ bit string. In this work, we denote the total number of processes (each node runs exactly one process) as $N$. Let $m_{p}$ denote the number of $\alpha$ bit strings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=x$. Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times y$ determinants. As each determinant is fully defined by its index through Eq. (2), our program does not require additional memory to store determinant composition information.
Figure 1: Distribution of the CI vector into segments across $N$ processes. Each process holds a strict subset of $m_{p}$ of the $\alpha$ strings. On a local process, the determinants are divided into segments (represented by vertical arrows), where a segment is composed of all determinants that share the same $\alpha$ string. For this figure, we define a displacement index $d_{p}=\sum_{i=0}^{p-1}m_{i}$ to show the offset of $\alpha$ strings on a given process.
2.2 Distributed Matrix-Vector Multiplication
In this work, we employ the traditional Davidson diagonalization method 17 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 18). However, both approaches store the CI vector entirely on a single node without exception. The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix is efficiently constructed using an MPI_ALLREDUCE operation. However, a key limitation is that the CI vector size is strictly constrained by the memory of a single node, posing a major bottleneck in utilizing supercomputing resources for larger determinant spaces.
In our program, the CI vector is distributed across nodes, and each node runs exactly one process. In this setting, each node $p$ only stores its corresponding CI vector segments. The computation of $\mathbf{W_{p}}$ is then given by:
$\mathbf{W_{p}}=\sum_{q}H\cdot\mathbf{U_{q}}.$ (3)
This implies that, in a single Davidson iteration, node $p$ theoretically needs to sequentially fetch data from all other nodes to compute $\mathbf{W_{p}}$. We define each individual fetch-and-compute operation (i.e., computing $\mathbf{W_{p}}\mathrel{+}=H\cdot\mathbf{U_{q}}$) as one step. Since there are $N$ computing nodes, each Davidson iteration consists of $N$ steps. Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix in memory, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
During the $N$ computation steps within a single Davidson iteration, each node operates independently without global synchronization. For example, once node $p$ completes the computation $\mathbf{W}_{p}\mathrel{+}=H\cdot\mathbf{U_{q}}$, it immediately fetches $\mathbf{U_{r}}$ from node $r$ and proceeds with $\mathbf{W_{p}}\mathrel{+}=H\cdot\mathbf{U_{r}}$. This design is crucial because computation times vary across different steps and nodes. Enforcing synchronization after each step would lead to unnecessary waiting and degrade computational efficiency. For data communication, we adopt the MPI_GET operation, a type of one-sided communication, allowing nodes to fetch data asynchronously without interrupting the computation on the target node, further improving efficiency.
2.3 Efficient Hamiltonian Computation for TPSCI
Due to the Slater-Condon rules, $H$ is a highly sparse matrix, so minimizing operations that involve zero elements is crucial for efficiency. In the DICE program, the nonzero Hamiltonian elements are precomputed and stored in a distributed manner along with their corresponding indices, allowing direct computations with these elements and avoiding operations on zeros.
While such precomputation strategies are effective, they require significant memory to store the nonzero elements. In contrast, our approach avoids explicit storage of Hamiltonian elements and instead exploits the inherent sparsity of $H$ to eliminate unnecessary computations. Here, we first analyze the structure of Hamiltonian matrix elements in SCI, which provides the foundation for our efficient computation strategy.
The Hamiltonian matrix element $\langle D_{w}|H|D_{u}\rangle$ can be expressed as $\langle S_{w}^{\alpha}|\langle S_{w}^{\beta}|H|S_{u}^{\alpha}\rangle|S_{u}^{\beta}\rangle$. According to the Slater-Condon rules, this matrix element can be nonzero only if:
$\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})+\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})\leq 2,$ (4)
where $\text{DIFF}(S_{w},S_{u})$ denotes the number of electron differences between bit strings $S_{w}$ and $S_{u}$. To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})$ and $b=\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})$. For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})=1$ and $\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})=1$. Throughout this work, we use the shorthand notations [2,0], [1,1], [1,0], [0,2], [0,1], and [0,0] to represent a total of six terms in which the matrix elements may be nonzero. The first five terms correspond to off-diagonal elements of the Hamiltonian matrix, which are always computed on-the-fly. The last term, [0,0], corresponds to diagonal elements, which, although also computable on-the-fly, are precomputed and stored in memory in our program to improve efficiency.
Data: p the current process, q the current step.
            SINGLE_LINK and DOUBLE_LINK lists.
1 for $S_{w}^{\alpha}$leftarrow $S^{\alpha}_{p,1}$ to $S^{\alpha}_{p,m_{p}}$ do
2      for $S_{u}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{u}^{\alpha}$)
4           if case = 2 then
5                for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
6                     $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
7                     Calculate [2,0] term
8
9                     end for
10
11                     else if case=1 then
12                          for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
13                               foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
14                                    Calculate [1,1] term
15
16                                    end foreach
17                                   $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
18                                    Calculate [1,0] term
19
20                                    end for
21
22                                    else if case=0 then
23                                         for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
24                                              foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
25                                                   Calculate [0,1] term
26
27                                                   end foreach
28                                                  $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
29                                                   Calculate [0,0] term
30
31                                                   end for
32
33                                                   end if
34
35                                                   end for
36                                                  $S_{u}^{\alpha}$leftarrow $S_{w}^{\alpha}$
37                                                   for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
38                                                        foreach $S_{u}^{\beta}$in the q-th batch of DOUBLE_LINK(:,$S_{w}^{\beta}$do
39                                                             Calculate [0,2] term
40
41                                                             end foreach
42
43                                                             end for
44
45                                                             end for
Algorithm 1 Computation algorithm for TPSCI on local process $p$ at step $q$. This algorithm describes how to compute $\langle D_{w}|H|D_{u}\rangle$, where $|D_{w}\rangle$ and $|D_{u}\rangle$ are determinants from processes $p$ and $q$, respectively.
Our novel algorithm for TPSCI is shown in Algorithm 1. Each Davidson iteration consists of $N$ steps. In step $q$, the local process $p$ performs the computation using data fetched from process $q$, which is detailed in Algorithm 1.
To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{w}^{\beta}\mid u=1,2,\dots,y\}$. These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{w}\rangle$ and $|D_{u}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency. This optimization is enabled due to the unique structure of TPSCI: for each $\alpha$ bit string, the associated $\beta$ bit string set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently.
As the DOUBLE_LINK array can be extremely large, it is distributed nearly evenly across all processes based on its indices. To keep track of this distribution, an auxiliary array is introduced to record the starting and ending indices of DOUBLE_LINK assigned to each process. Consequently, the computation of the [0,2] term is performed batch by batch over the $N$ steps (as shown in Lines 27-32 of Algorithm 1). Since the calculation of the [0,2] term is performed only in the loop over $S_{w}^{\alpha}$, each process $p$ will fetch data from the target process $q$ for the DOUBLE_LINK only once per Davidson step.
Roughly speaking, the computational cost of TPSCI scales as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, where $N_{\mathrm{SCI}}$ is the number of selected determinants, $N_{\mathrm{occ}}$ the number of occupied orbitals, and $N_{\mathrm{vir}}$ the number of virtual orbitals. This can be further reduced by considering the sparsity of excitation links. The sparsity of single excitation links roughly scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, in which $N_{\mathrm{FCI}}$ is the FCI dimension, and thus the overall scaling is more precisely given by $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$.
2.4 Strategies for Efficient Parallel Execution in TPSCI
As described above, we distribute the CI vector across nodes to overcome single-node memory limitations, enabling computations to fully utilize the total available memory. However, this distributed storage introduces new challenges: a computation that could previously be completed in a single step now requires $N$ steps, with each step requiring data retrieval from a target process [as shown in Eq. (3)]. As the number of nodes increases, the overall MPI communication volume seemingly scales as $N_{\text{SCI}}^{2}$, making communication latency a potential bottleneck for computational efficiency. To address this issue, we have developed a series of strategies to reduce communication overhead and improve parallel execution.
2.4.1 Avoiding Unnecessary MPI Data Transfers
In our previous analysis, we mentioned that each local process $p$ needs to fetch data from other processes. However, not all data from the target process is required. Specifically, if all $\alpha$ bit strings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{u})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{u}$.
In our implementation, such information is precomputed and stored, allowing each process to identify and skip unnecessary data transfers, thereby significantly improving communication efficiency. As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot y$. It can be easily shown that the computational cost and the data communication volume on each process share the same scaling. Consequently, the overall MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$ rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TPSCI on supercomputers.
2.4.2 Minimizing Long-Distance Communication in MPI
Each node in the supercomputer is assigned a unique node identity number (ID) ranging from 00 to $N-1$. When suitably ordered, messages sent between nodes with closer IDs will on average require fewer hops compared to those with larger ID differences.
Based on this, we sort the $\alpha$ bit strings according to their excitation levels relative to the Hartree-Fock (HF) $\alpha$ bit string, and then assign them to nodes in increasing order. Hence, lower-excitation $\alpha$ bit strings are assigned to lower-numbered nodes, whereas higher-excitation strings are assigned to higher-numbered nodes.
As a result, communication between distant nodes is significantly reduced because the excitation differences between their assigned $\alpha$ bit strings are more likely to be all greater than 2, making communication unnecessary. This strategy is crucial for reducing MPI communication congestion, a well-known challenge in large-scale high-performance computing, by ensuring that each node primarily fetches data from nearby nodes, thereby improving parallel efficiency.
2.4.3 Achieving Computational Load Balance in MPI
In MPI-based parallel computing, maintaining computational load balance is crucial because the slowest process determines the overall execution time. The challenge is how to estimate each process’s workload in advance. As shown in Algorithm 1, the most time-consuming computation is the evaluation of the [1,1] term. Therefore, the number of single excitations associated with each $\alpha$ bit string can be used as a rough estimate of the computational workload for its corresponding segment. Consequently, the total computational workload of a process can be estimated by summing the single excitation counts of all $\alpha$ bit strings assigned to it.
As described above, we assign the $\alpha$ bit strings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bit string. This ordering is crucial in our TPSCI program to minimize long-distance communication between nodes. However, balancing memory usage and computational workload simultaneously presents a challenge. Since the total number of $\alpha$ bit strings is fixed, the average number per node is also predetermined. Similarly, the total number of single excitation counts of all $\alpha$ bit strings is fixed, meaning that the average single excitation count per process is also predetermined.
To quantify the balance between memory and computation, we define two expansion factors: the memory expansion factor, which is the ratio of the assigned $\alpha$ bit string count to the average, and the computation expansion factor, which is the ratio of the assigned single excitation counts to the average. Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can lead to severe computational imbalance, causing some nodes to finish significantly earlier and remain idle while waiting for others. Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may result in a highly uneven memory distribution, with some nodes exceeding their memory limits, forcing an increase in the total number of nodes and leading to inefficient resource usage.
To address this issue, we adopt a balanced strategy where both expansion factors are kept moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive.
2.4.4 Overlapping Computation and Data Transfer to Reduce MPI Latency
As mentioned above, each node in our program runs a single process. Modern computers are typically multi-core; for example, each node on supercomputer Fugaku has 48 CPU cores. Even though we minimize unnecessary data transfers and reduce long-distance communication, some degree of MPI communication congestion remains unavoidable. The question then arises: how can we further mitigate its impact on the overall computation wall time?
To mitigate this issue, one core is assigned to perform MPI data transfers, fetching the required data for the next step, while the remaining 47 cores perform parallel computation for the current step using OpenMP. Let $t_{\text{cal}}$ and $t_{\text{data}}$ represent the computation time and the data retrieval time of one step, respectively. If $t_{\text{data}}>t_{\text{cal}}$, we define the delay time to the wall time as $t_{\text{delay}}=t_{\text{data}}-t_{\text{cal}}$. Conversely, if $t_{\text{data}}\leq t_{\text{cal}}$, then $t_{\text{delay}}=0$, since MPI communication congestion during data transfer—if it occurs—does not prolong the wall time. The total delay time for a Davidson iteration, denoted as $T_{\text{delay}}$, is given by the sum of $t_{\text{delay}}$ over all $N$ steps.
2.4.5 “Check if Busy” Strategy for Communication
During data fetching from target processes, we employ a strategy called “Check if Busy”. Specifically, when process A uses MPI_GET to fetch data from process B, both A and B are marked as BUSY. During this period, no other process can fetch data from either A or B; instead, they will first fetch data from processes that are NOT BUSY. This strategy helps prevent multiple processes from simultaneously accessing the same process, thereby reducing communication congestion and improving overall parallel efficiency.
3 Results
Although the TPSCI program in this work is designed for “QC+HPC”, this study is an independent effort and does not yet incorporate QC-based bit string selection. Instead, we perform FCI calculations to provide widely applicable benchmark data. Consequently, the computational scaling follows $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, which is higher than the $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$ scaling of the Handy-Knowles FCI algorithm. Our goal is not to compete with existing FCI methods but to showcase the effectiveness of the newly designed TPSCI algorithm and its high efficiency in parallel execution.
In actual calculations, we further introduced molecular symmetry to reduce both memory usage and computational cost. All HF reference states and their corresponding FCIDUMP files were generated using PySCF. 19 For double-precision TPSCI, convergence was defined as an energy difference of less than $1\times 10^{-7}$ Hartree between consecutive iterations. For single-precision TPSCI, due to fewer significant digits, a looser criterion of less than $1\times 10^{-6}$ Hartree was used. Notably, the final total energies obtained from both versions differ only in the sixth decimal place, indicating that the single-precision SCI calculation is accurate to $0.01$ milliHartree (mH), which is sufficient for chemical applications.
We perform experiments on supercomputer Fugaku, which is a CPU only architecture based on the A64FX processor. Each node has 48 cores, and only 32 GiB of memory, making processing and memory distribution essential. The code was compiled with the Fujitsu compilers using the -Kfast,parallel,openmp,SVE options.
3.1 N2 with the aug-cc-pVDZ Basis Set
Using $D_{2h}$ symmetry with frozen core orbitals, the N2 molecule with the aug-cc-pVDZ basis set has $1.47\times 10^{11}$ determinants. We first evaluate the computational time of a single Davidson iteration using different numbers of nodes. As shown in Table 1, the program’s wall time, defined as the wall time of the slowest process, remains within 116% of the average wall time across all processes, demonstrating good MPI load balancing.
Table 1: Parallel performance of TPSCI for N2 computed with the aug-cc-pVDZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.
Nodes Avg. Max. $T_{\text{delay}}$ NHs
480 2259 2484 115 331.2
960 1179 1292 7 344.5
1440 834 917 5 366.8
1920 655 740 14 394.7
2400 554 618 11 412.0
2880 483 552 28 441.6
A particularly notable result is that the total delay time ($T_{\text{delay}}$) does not increase significantly as the number of nodes grows, demonstrating strong parallel efficiency. Generally, increasing the number of nodes worsens communication congestion, which in turn increases MPI latency. In fact, before implementing the optimization strategies introduced in Section 2.4, the measured $T_{\text{delay}}$ values exhibited a significant growth trend with increasing nodes, sometimes even resulting in longer wall times despite using more nodes. However, after applying these strategies, $T_{\text{delay}}$ was significantly reduced, allowing the wall time to decrease consistently with increasing nodes. Although the total node-hours (NHs) still increase with node count (i.e. decreasing parallel efficiency), the substantial wall time reduction confirms the effectiveness of our parallelization strategies.
Table 2: Potential Energy Surface of N2 in Hartree computed with the aug-cc-pVDZ basis set for different bond lengths (R). We report the HF energy and the FCI energies computed with double (FP64) and single (FP32) precision.
$R$ (Bohr) HF FP64 FP32
2.118 -108.956105 -109.297336 -109.297335
2.4 -108.875323 -109.258928 -109.258927
2.7 -108.748038 -109.182160 -109.182160
3.0 -108.618929 -109.108790 -109.108789
3.6 -108.401280 -109.016333 -109.016335
4.2 -108.242916 -109.986038 -109.986044
The FCI energies of the potential energy surface of N2 are shown in Table 2. The double-precision and single-precision results are nearly identical, with the maximum difference being only 0.006 mH at $R=4.2$ Bohr. This data validates the accuracy and reliability of our single-precision implementation for TPSCI calculations.
3.2 CN with the cc-pVTZ Basis Set
As our program is capable of performing open-shell system calculations, we also provide benchmark data for CN, computed with C2v symmetry, frozen core orbitals, and a bond length of 1.16 Å. For the cc-pVDZ and cc-pVTZ basis sets, these calculations require $2.46\times 10^{8}$ and $4.86\times 10^{11}$ determinants, respectively. We emphasize here that naively storing (in single precision) even a single CI vector alone on each node would require about 1.8 TB of memory; however, using our distributed memory scheme, we can unlock the full memory of supercomputer Fugaku and perform such a large calculation. We also report the strong scaling performance of this calculation in Table 3.
Table 3: Parallel performance of TPSCI for CN computed with the cc-pVTZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.
Nodes Avg. Max. $T_{\text{delay}}$ NHs
2400 4019 4400 401 2933
4800 2231 2534 490 3379
7200 1497 1681 242 3362
In Table 4, we present a comparison of the energy values computed by FCI with various orders of coupled cluster theory computed using NWChem 20. The FCI values represent an improvement of approximately 1-2 mH over CCSDT 21, 22, 23. For a system of this size, CCSDTQ 24, 25 provides an excellent approximation, with an error of only about 0.2 mH. In addition, as our calculations represent one of only a few systems that can be computed with a triple-$\zeta$ quality basis set by FCI, we also perform $2/3$ extrapolation of this data to the complete basis set limit (CBS). We extrapolate the HF energy ($E_{c}^{\text{HF}}$) of a given cardinality ($c$) as CBS($l$/$m$) $=E^{\text{HF}}_{m}+\frac{(E^{\text{HF}}_{m}-E^{\text{HF}}_{l})m^{-5}}{m^{-5}-l^{-5}}$ and the correlation energy ($E_{c}^{\text{Cor}}$) as CBS($l$/$m$) $=E^{Cor}_{m}+\frac{(E^{\text{Cor}}_{m}-E^{\text{Cor}}_{l})m^{-3}}{m^{-3}-l^{-3}}$ 26. We hope that this data for a challenging system can be used for benchmarking new methods.
Table 4: Energy values of CN using different methods and basis sets. We denote the two point extrapolation using basis sets of cardinality $l$ and $m$ as CBS($l$/$m$).
Basis Set HF CCSD CCSD(T) CCSDT CCSDTQ FCI
cc-pVDZ -92.197616 -92.475781 -92.489728 -92.490269 -92.491601 -92.491812
cc-pVTZ -92.219490 -92.547372 -92.567450 -92.567478 -92.569054 -92.569252
CBS(2/3) -92.222807 -92.571623 -92.594283 -92.594095 -92.592456 -92.595966
4 Conclusion
Inspired by the idea of TPSCI proposed by an IBM team in 2024, we proposed and implemented a novel and efficient algorithm in this work. To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations. Additionally, to further reduce memory usage, we store only the diagonal elements of the Hamiltonian matrix while computing the off-diagonal elements on-the-fly. As a result, we successfully handled up to $4.86\times 10^{11}$ determinants, surpassing the previously reported largest SCI calculation ($2\times 10^{9}$ determinants) by more than two orders of magnitude.
Meanwhile, large-scale parallel computations on supercomputers face a critical challenge: increasing the number of nodes often leads to severe communication congestion, significantly degrading parallel efficiency. To address this, we have developed and implemented a series of optimization strategies tailored to TPSCI, including eliminating unnecessary communication, reducing long-distance data transfers, maintaining balanced computational load and memory distribution in MPI, hiding communication by prefetching, and detecting target process statuses before data transfer. The results demonstrate that even with up to 7,200 nodes, the program’s wall time remains within 116% of the average across all processes, and communication delays ($T_{\text{delay}}$) do not increase significantly with the number of nodes. Such exceptional performance lays a solid foundation for the development of TPSCI in the future.
Since this study is an independent effort of our research group, it does not yet incorporate determinant selection using QC. Therefore, we directly performed FCI calculations on N2 with the aug-cc-pVDZ basis set and CN with the cc-pVTZ basis set, both of which are larger than the cc-pVDZ basis set reported in previous FCI studies of these molecules. These data can be used as benchmark in future quantum chemistry studies.
Looking ahead, the efficient distributed and parallelized framework developed in this work provides a strong foundation for integrating TPSCI with QC approaches. In particular, quantum computers could be leveraged to identify important bit strings, which could then be used to construct a spin-adapted determinant space via tensor products. These determinants would be efficiently processed using our highly efficient and scalable program. The combination of quantum bit string selection with large-scale classical computation has the potential to greatly enhance electronic structure calculations, enabling the accurate simulation of even larger and more complex molecular systems.
Acknowledgments
The authors thank Muneaki Kamiya for suggesting a restart strategy in the Davidson algorithm to reduce memory usage, and Himari Pathak for discussions on applications.